OpenStudy (anonymous):

[EDIT: Work down below, appreciate if anyone could look at it] What is the reciprocal of 3 - √-5? My teacher may have skipped over this part in note-taking, so I'm at a stalemate trying to figure out what to do.

Nnesha (nnesha):

reciprocal of a/b is b/a just flip the fraction reciprocal of 2(which is same as 2/1) is 1/2

OpenStudy (anonymous):

I figured out that part, but I accidentally missed out on the answer choices. 1) 3/2 + 3/2i 2) 3/2 - 3/2i 3) 3/4 + i√5/4 4) 3/4 - i√5/4 is there more to it than simply the reciprocal? That's what I believe

Nnesha (nnesha):

alright so you have to simplify that what's the reciprocal of `3-sqrt{-5}` ??

OpenStudy (anonymous):

1/3 - √-5

Nnesha (nnesha):

right \[\rm \frac{ 1 }{ 3-\sqrt{-5} }\] multiply both numerator and denominator by the conjugate

Nnesha (nnesha):

do you know what's conjugate ?

Nnesha (nnesha):

do you know what's conjugate ?

OpenStudy (anonymous):

conjugate means change the sign, so \[\frac{ 1 }{ 3 - \sqrt{-5} } * \frac{ 3 + \sqrt{-5} }{ 3 + \sqrt{-5} } = \frac{ 3 + \sqrt{-5} }{ 9 + 5 } = \frac{ 3 + 5i }{ 14 }\] I feel as if I did something wrong though

Nnesha (nnesha):

hmm 5 would stay under the radical sign

Nnesha (nnesha):

\[\sqrt{-1} \rightarrow \sqrt{-1} \sqrt{5} \rightarrow i \sqrt{5}\]

OpenStudy (anonymous):

http://imgur.com/Qp01FiP If it's too small, let me know and I'll type it up

Nnesha (nnesha):

hmm

Nnesha (nnesha):

your work is correct. check your answer choices are you sure its not 14 instead 4 ?